THE VISIBLE SPECTRUM OF ZIRCONIUM DIOXIDE, ZrO_{2}

ANH LE AND \underline{TIMOTHY C. STEIMLE}, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287; VARUN GUPTA AND JOHN P. MAIER , Department of Chemistry, University of Basel, Basel, Switzerland.

Zirconia (ZrO_{2}) is an important material with applications in microelectronics, catalysis, and ceramics. Previously the photoelectron spectrum^{a}^{,}^{b} of the ZrO_{2} anion and the pure rotation spectrum of the neutral^{c} have been recorded and analyzed. Here we present the first observation of the visible spectrum of ZrO_{2} via laser induced fluorescence (LIF) and resonance enhanced multi-photon ionization (REMPI). The LIF spectrum was recorded between 17000-18900 cm^{−1} at a resolution of 0.2 cm^{−1} using pulsed dye laser excitation and tentatively analyzed to give harmonic vibrational parameters ω_{1}, ω_{2}, and ω_{3} for the Ã^{1}B_{2} state of 495(1) cm^{−1}, 150(3) cm^{−1} and 1045(4) cm^{−1}, respectively. Dispersed fluorescence spectra of thirteen bands were recorded and analyzed to give harmonic vibrational parameters ω_{1}, ω_{2}, and ω_{3} for the X^{1}A_{1} state of 898 (1) cm^{−1}, 287(3) cm^{−1} and 808(4) cm^{−1}, respectively. The radiative lifetimes of numerous bands have been measured and analyzed. Franck-Condon factors were calculated and used to model the REMPI and excitation spectra. A comparison with TiO_{2} is made^{d}.

aO.C. Thomas; S. Xu; T.P. Lippa; and K.H. Bowen Journal of Cluster Science 10(4) 525 ,1999.

bW.Zheng; K.H. Bowen JPCA. 109 11521, 2005

cD.J. Brugh; R.D. Suenram JCP 11(8), 3526,1999

dX. Zhuang; A. Le; T.C. Steimle; R. Magarajan; V. Gupta; and J.P. Maier PCCP 12 15018, 2010